# Quantitative imaging of nonlinear spin-wave propagation using diamond quantum sensors

### v1

@@ III.3 Nonlinear behavior of spin-wave amplitude
Now, we conduct a quantitative analysis of the spin-wave amplitude in the nonlinear regime. We convert the obtained microwave amplitude distribution from spin waves $B_{\mathrm{sw}}(x)$ into a spin-wave amplitude distribution. $B_{\mathrm{sw}}(x)$ is related to the spin-wave amplitude distribution in the $x$-direction $m_{x}(x)$ by the following equation [21, 22, 26]:
$B_{\mathrm{sw}}(x)=\frac{\mu_{0}}{2}e^{-k_{x}(x)z_{d}}(1-e^{-k_{x}(x)d})(1+\eta_{k_{x}(x)})m_{x}(x),\$ (5)
where $\eta_{k_{x}(x)}$ is the ellipticity of the spin-wave precession, which can be obtained analytically from the experimental condition (See Appendix B of Ref. [26] for details). Figure 3(a) shows the results of converting $B_{\mathrm{sw}}(x)$, shown in Fig. 2(a), into $m_{x}(x)$ using Eq. (5). As with the microwave amplitude distribution, the decay of the spin-wave amplitude near the stripline becomes more pronounced above $P_{\mathrm{mw}}=19\,\mathrm{dBm}$. The nonlinearity of the spin-wave amplitude is also reproduced in numerical simulations (see Supplemental Materials for details), which validates our results.
Figure 3(b) displays a heatmap of $m_{x}(x)$ from Fig. 3(a) as a function of the input microwave amplitude $H_{\mathrm{mw}}\,(\coloneqq\sqrt{P_{\mathrm{mw}}})$. Up to approximately $H_{\mathrm{mw}}\sim 10\,\sqrt{\mathrm{mW}}$, the spin-wave amplitude around the left edge of the field of view (small $x$ region) increases with the increase of $H_{\mathrm{mw}}$, and gradually decays along the propagation direction. Beyond this region, however, the amplitude does not increase and decays rapidly from the left edge, signaling peculiar nonlinearity.
Figure 3: (a) Calculated spin-wave amplitude distribution $m_{x}(x)$ for each microwave input power for YIG-A. (b) Heatmap of the spin-wave amplitude distribution as a function of the microwave input power. (c)(d) Dependence of the spin-wave amplitude on input microwave amplitude (c) at $x=4\,\mathrm{\mu m}$ and (d) at $x=13\,\mathrm{\mu m}$.
Figure 3(c) shows the spin-wave amplitude at $x=4\,\mathrm{\mu m}$ as a function of the input microwave amplitude. The red dashed line represents a linear fit through the origin based on data from the two smallest microwave input amplitudes ($P_{\mathrm{mw}}=5,11\,\mathrm{dBm}$). The spin-wave amplitude begins to deviate from linear behavior around $\mu_{0}m_{x}=20\,\mathrm{mT}$, increases until approximately $30\,\mathrm{mT}$, and then begins to decay. This deviation from linearity at a certain microwave power is consistent with previous studies using VNA and BLS [12, 8, 9, 13]. Figure 3(d) shows the dependence of the spin-wave amplitude on $x=13\,\mathrm{\mu m}$. The amplitude deviates from linear behavior at the same $H_{\mathrm{mw}}$ as $x=4\,\mathrm{\mu m}$, saturates around $20\,\mathrm{mT}$, and then begins to decay.
@@ III.4 Theoretical interpretation: four-magnon scattering
The observed nonlinear decay of the spin-wave amplitude can be attributed to multi-magnon scattering. In this process, excited magnons characterized by frequency $f_{\mathrm{sw}}$ and wavenumber $k$ are scattered into magnons with different frequencies and wavenumbers while conserving the total energy and wavenumber, and this scattering process becomes pronounced beyond a certain spin wave amplitude threshold. In this subsection, according to the literature [36, 11], we first formulate the time evolution of magnon operators governed by a Hamiltonian that includes nonlinear terms. We calculate the threshold magnon amplitude at which nonlinearity becomes significant. We then perform a quantitative comparison with experimental results.
@@
where the first term represents the linear term describing the dispersion relation of the spin waves, while the second term corresponds to the nonlinear four-magnon scattering process characterized by the coupling strength $H_{\bm{k_{1}},\bm{k_{2}},\bm{k_{3}},\bm{k_{4}}}$ (see Supplemental Materials for the specific expressions for each term and their derivations). Although nonlinear processes such as three-magnon scattering also appear among the nonlinear terms, they do not contribute due to conservation laws governing frequency and wavenumber. Therefore, in this case, we only consider the four-magnon scattering term described above (see Supplemental Materials for details).
@@
Here, we consider the situation where only magnons with wavenumber $\bm{k}$ are excited by a microwave field. These excited magnons subsequently undergo four-magnon scattering such that $\bm{k_{1}}=\bm{k}+\bm{q}$, $\bm{k_{2}}=\bm{k}-\bm{q}$, and $\bm{k_{3}}=\bm{k_{4}}=\bm{k}$ in Eq. (6) [36, 11]. In the rotating frame defined as $\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}\,(=\hat{c}_{\bm{k}-\bm{q}}e^{i\omega_{\bm{k}-\bm{q}}t})$, the time evolution of the magnon annihilation operator obeys the following differential equation:
@@
where we introduce the decay term $\Gamma_{\bm{k}-\bm{q}}$. Following the same discussion, the time evolution of $\hat{c}^{{}^{\prime}{\dagger}}_{\bm{k}+\bm{q}}$ obeys the following differential equation:
@@
For the nonlinear term in Eqs. (8) and (9), we assume that $\hat{c}^{{}^{\prime}}_{\bm{k}}$ attains the steady-state value $\expectationvalue{\hat{c}^{{}^{\prime}}_{\bm{k}}}$ by resonant microwave excitation. By combining Eqs. (8) and (9), we derive the following differential equation for $\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}$:
@@
Now, we define the magnon generation rate $\gamma_{\bm{k}-\bm{q}}$ by assuming that $\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}\propto e^{\gamma_{\bm{k}-\bm{q}}t}$. Taking the limit as $\bm{q}\rightarrow 0$ in Eq. (10), we can derive
@@
This equation indicates that the generation rate of magnons produced by scattering changes from negative to positive when the magnon amplitude increases from 0. The resulting sign change corresponds to a threshold at which the excited coherent spin waves begin to decay via four-magnon scattering. At this threshold, the magnon amplitude is given by
@@
where $u_{\bm{k}}$ and $v_{\bm{k}}$ are dimensionless parameters of the Bogoliubov transformation [37] (see Supplemental Materials for details), and $M_{\mathrm{s}}$ is the saturation magnetization of the magnetic film.
We calculate the spin-wave amplitude threshold corresponding to the experimental results using this expression. Table 1 shows the relevant experimental parameters. For YIG-A, we use the parameters listed in the left column of Table 1. The saturation magnetization is taken as $M_{\mathrm{s}}=141\,\mathrm{mT}$, as measured by SQUID (superconducting quantum interference device) magnetometry at 300 K. The anisotropy field $H_{\mathrm{ani}}$ is estimated from the difference between the effective magnetization $M_{\mathrm{eff}}=169.6\,\mathrm{mT}$ and the saturation magnetization by $H_{\mathrm{ani}}=M_{s}-M_{\mathrm{eff}}$. The decay rate $\Gamma_{\bm{k}}$ is calculated from the spin-wave lifetime, which is obtained by dividing the decay length $l_{d}=54.21\,\mathrm{\mu m}$ (determined from fitting the data at $P_{\mathrm{mw}}=11\,\mathrm{dBm}$) by the group velocity of the spin waves. The four-magnon scattering coefficient $\lambda_{\bm{k}}$ and the Bogoliubov transformation parameters $u_{\bm{k}}$ and $v_{\bm{k}}$ are calculated by the analytical expressions using the experimental parameters (see Supplemental Materials for the expressions). Based on these parameters, the spin-wave amplitude threshold for the instability due to four-magnon scattering is calculated to be $\mu_{0}m^{\mathrm{th}}_{x,\bm{k}}=17.8\,\mathrm{mT}$.
It is significant to examine how the above discussion applies to our experiment. The black dotted lines in Figs. 3(c) and (d) indicate the threshold amplitude thus obtained. In particular, regarding Fig. 3(c), which displays the spin-wave amplitude at $x=4\,\mathrm{\mu m}$ near the stripline, the threshold for the emergence of nonlinearity and the experimental results agree satisfactorily. This result demonstrates that the nonlinear dynamics can be quantitatively visualized using NV centers. The ability to experimentally detect and quantitatively characterize the transition from linear to nonlinear regimes, consistent with theoretical predictions, constitutes the central achievement of the present work.
@@
External magnetic field $B_{0}$ $25.17\,\mathrm{mT}$ $22.45\,\mathrm{mT}$
@@
Anisotropy field $H_{\mathrm{ani}}$ $-28.6\,\mathrm{mT}$ $-15.6\,\mathrm{mT}$
Spin-wave frequency $f_{\mathrm{sw}}$ $2166\,\mathrm{MHz}$ $2243.1\,\mathrm{MHz}$
@@
Four-magnon scattering coefficient $\lambda_{\bm{k}}$0.155 -0.173
Bogoliubov transformation parameter $u_{\bm{k}}$ 1.11 1.10
@@ III.5 Magnetization reduction in the nonlinear regime
Having examined the spin-wave amplitude, we next analyze the spatial distribution of the wavenumbers. The observed increase in the wavenumber reflects a decrease in the static magnetization along the $y$-direction (i.e., in the direction of the saturation magnetization) $M_{y,s}$ and corresponding effective magnetization $M_{\mathrm{eff}}$ due to an increase in the spin-wave amplitude [Fig.4(a)] [13, 14]. To calculate the effective magnetization, we utilize the dispersion relation of surface spin waves [Eq. (1)]. At each position, the effective magnetization $M_{\mathrm{eff}}(x)$ corresponding to the spin-wave frequency $f_{\mathrm{sw}}=2165.8\,\mathrm{MHz}$ is calculated from the wavenumber $k_{x}(x)$. As the dispersion relation holds for a small spin-wave amplitude, there may be some modulation of the dispersion relation from Eq. (1) for a large spin-wave amplitude, particularly in nonlinear regions. However, the coherent spin-wave amplitude with the excitation microwave frequency decays rapidly near the stripline in the nonlinear region, and the amplitude is small compared to the saturation magnetization across most areas. Therefore, we utilize the dispersion relation of the surface spin waves in Eq. (1) for the analysis approximately.
Figure 4(c) plots the resulting spatial distribution of the static magnetization $M_{y,\mathrm{s}}(x)$. We convert the effective magnetization distribution $M_{\mathrm{eff}}(x)$ into the static magnetization distribution $M_{\mathrm{y,s}}(x)$ by removing the contribution from the surface anisotropy $\quantity(M_{\mathrm{y,s}}(x)=M_{\mathrm{eff}}(x)+H_{\mathrm{ani}})$. Figure 4(c) shows that, in regions with small microwave input power, though there is uniform resuction as the power increases, the static magnetization remains almost constant within the field of view. However, in addition to the uniform reduction, significant variation within the field of view appears above $P_{\mathrm{mw}}=19\,\mathrm{dBm}$, and at the maximum input power ($P_{\mathrm{mw}}=25\,\mathrm{dBm}$), a variation of about $6\,\mathrm{mT}$ within the field of view is observed.
@@
It is interesting to examine how the surface spin waves contribute to the reduction in static magnetization [Fig. 4(b)]. When considering only the contributions from the coherent surface spin waves with the excitation frequency, the magnetization in the $y$-direction at a specific position and time is expressed as
$\begin{split}M_{y}(t)&=\sqrt{M^{2}_{s}-m^{2}_{x}(t)-m^{2}_{z}(t)},\\ \end{split}$ (15)
@@
[…]
Figure 4(d) shows the effective static magnetization distribution $M^{{}^{\prime}}_{y,\mathrm{s}}(x)$, which is calculated from the spin-wave amplitude distribution in Fig. 3(a) using Eq. (17). The left edge of the plot is defined at $x=0.91\,\mathrm{\mu m}$. At low microwave input powers in the linear region, $M^{{}^{\prime}}_{y,\mathrm{s}}$ shows a slight decrease only at the left edge of the field of view ($x\lesssim 10$ $\mathrm{\mu m}$), which is in clear contrast to the observation that the static magnetization derived from the wavenumber [$M_{\mathrm{y,s}}(x)$ shown in Fig. 4(c)] exhibits a prominent decrease across the entire field of view. These results imply that the decay of the surface spin waves is not solely due to relaxation into the lattice but also involves scattering by other magnons, such as thermal magnons. In other words, the excited surface spin waves are transformed into other modes, leading to a persistent reduction in static magnetization across the field of view [37].
Around maximum microwave input power in the nonlinear regime, the reduction in static magnetization attributable to surface spin waves becomes almost negligible at a distance of about $5\,\mathrm{\mu m}$ from the stripline. In contrast, the static magnetization estimated from the wavenumber [Fig. 4(c)] continues to change gradually across the field of view. This observation suggests that most of the modulation in magnetization is caused by additional magnons generated through four-magnon scattering from the excited surface spin waves.
As a note, we investigate the possible role of Joule heating resulting from the application of a large-amplitude microwave, which might cause the sample to heat up. Such heating could potentially create a temperature gradient along the $x$-direction, thereby causing a spatial variation in the saturation magnetization and wavenumber. However, we can safely exclude this possibility. We perform similar measurements at a different magnetic field, where the spin-wave excitation is more inefficient and the spin waves are less likely to reach the nonlinear region, and confirm that the wavenumber along the $x$-direction is almost constant even at high microwave power (see Supplemental Materials for details).

### v2

@@ III.3 Nonlinear behavior of spin-wave amplitude
Now, we conduct a quantitative analysis of the spin-wave amplitude in the nonlinear regime. We convert the obtained microwave amplitude distribution from spin waves $B_{\mathrm{sw}}(x)$ into the corresponding spin-wave amplitude distribution. $B_{\mathrm{sw}}(x)$ is related to the spin-wave amplitude distribution in the $x$-direction $m_{x}(x)$ by the following equation [21, 22, 26]:
$B_{\mathrm{sw}}(x)=\frac{\mu_{0}}{2}e^{-k_{x}(x)z_{d}}(1-e^{-k_{x}(x)d})(1+\eta_{k_{x}(x)})m_{x}(x),\\$ (5)
where $\eta_{k_{x}(x)}$ is the ellipticity of the spin-wave precession, which can be obtained analytically from the experimental conditions (See Appendix B of Ref. [26] for details). Figure 3(a) shows the results of converting $B_{\mathrm{sw}}(x)$, shown in Fig. 2(a), into $m_{x}(x)$ using Eq. (5). As with the microwave amplitude distribution, the decay of the spin-wave amplitude near the stripline becomes more pronounced above $P_{\mathrm{mw}}=19\,\mathrm{dBm}$. The nonlinearity of the spin-wave amplitude is also reproduced by numerical simulations (see Supplemental Materials for details [30]), which validates our results.
Figure 3(b) displays a heatmap of $m_{x}(x)$ from Fig. 3(a) as a function of the input microwave amplitude $H_{\mathrm{mw}}\,(\coloneqq\sqrt{P_{\mathrm{mw}}})$. Up to approximately $H_{\mathrm{mw}}\sim 10\,\sqrt{\mathrm{mW}}$, the spin-wave amplitude around the left edge of the field of view (small $x$ region) increases with the increase of $H_{\mathrm{mw}}$, and gradually decays along the propagation direction. Beyond this range, however, the amplitude no longer increases and instead decays rapidly from the left edge, signaling peculiar nonlinearity.
Figure 3: (a) Calculated spin-wave amplitude distribution $m_{x}(x)$ for each input microwave power for YIG-A. (b) Heatmap of the spin-wave amplitude distribution as a function of the input microwave amplitude. (c)(d) Dependence of the spin-wave amplitude on input microwave amplitude (c) at $x=4\,\mathrm{\mu m}$ and (d) at $x=13\,\mathrm{\mu m}$.
Figure 3(c) shows the spin-wave amplitude at $x=4\,\mathrm{\mu m}$ as a function of the input microwave amplitude. The red dashed line represents a linear fit through the origin based on data from the two smallest input powers ($P_{\mathrm{mw}}=5,11\,\mathrm{dBm}$). The spin-wave amplitude begins to deviate from linear behavior around $\mu_{0}m_{x}=20\,\mathrm{mT}$, increases up to approximately $30\,\mathrm{mT}$, and then begins to decay. This deviation from linearity at a certain microwave power is consistent with previous studies using VNA and BLS [12, 8, 9, 13]. Figure 3(d) shows the dependence of the spin-wave amplitude at $x=13\,\mathrm{\mu m}$. The amplitude deviates from linear behavior at the same $H_{\mathrm{mw}}$ as $x=4\,\mathrm{\mu m}$, saturates around $20\,\mathrm{mT}$, and then begins to decay.
@@ III.4 Theoretical interpretation: four-magnon scattering
The observed nonlinear decay of the spin-wave amplitude can be attributed to multi-magnon scattering. In this process, excited magnons characterized by frequency $f_{\mathrm{sw}}$ and wavevector $\bm{k}$ are scattered into magnons with different frequencies and wavevectors while conserving the total energy and wavevector, and this scattering process becomes pronounced once the spin-wave amplitude exceeds a certain threshold. In this subsection, following Refs. [39, 40, 41, 42, 11, 2, 15], we first formulate the time evolution of magnon operators under the Hamiltonian that includes nonlinear terms and calculate the threshold magnon amplitude at which nonlinearity becomes significant. We then perform a quantitative comparison with the experimental results.
@@
where the first term represents the linear term describing the dispersion relation of the spin waves, while the second term corresponds to the nonlinear four-magnon scattering process characterized by the coupling strength $H_{\bm{k_{1}},\bm{k_{2}},\bm{k_{3}},\bm{k_{4}}}$ (see Supplemental Materials for the explicit expressions for each term and their derivations [30]). Although nonlinear processes such as three-magnon scattering also appear among the nonlinear terms, they do not contribute due to conservation laws governing frequency and wavevector. Therefore, we consider only the four-magnon scattering term described above (see Supplemental Materials for details [30]).
@@
Here, we consider the situation where only magnons with wavevector $\bm{k}$ are excited by a microwave field. These excited magnons subsequently undergo four-magnon scattering such that $\bm{k_{1}}=\bm{k}+\bm{q}$, $\bm{k_{2}}=\bm{k}-\bm{q}$, and $\bm{k_{3}}=\bm{k_{4}}=\bm{k}$ in Eq. (6) [42, 11]. In the rotating frame defined as $\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}\,(=\hat{c}_{\bm{k}-\bm{q}}e^{i\omega_{\bm{k}-\bm{q}}t})$, the time evolution of the magnon annihilation operator obeys the following differential equation:
@@
where we have introduced the decay term $\Gamma_{\bm{k}-\bm{q}}$. Following the same discussion, the time evolution of $\hat{c}^{{}^{\prime}{\dagger}}_{\bm{k}+\bm{q}}$ obeys the following differential equation:
@@
For the nonlinear term in Eqs. (8) and (9), we assume that $\hat{c}^{{}^{\prime}}_{\bm{k}}$ attains the steady-state value $\expectationvalue{\hat{c}^{{}^{\prime}}_{\bm{k}}}$ under resonant microwave excitation [43]. By combining Eqs. (8) and (9), we derive the following differential equation for $\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}$:
@@
Now, we define the magnon generation rate $\gamma_{\bm{k}-\bm{q}}$ by assuming that $\hat{c}^{{}^{\prime}}_{\bm{k}-\bm{q}}\propto e^{\gamma_{\bm{k}-\bm{q}}t}$. Taking the limit $\bm{q}\rightarrow 0$ in Eq. (10), we obtain
@@
This equation indicates that the generation rate of magnons produced by scattering changes from negative to positive as the magnon amplitude increases from zero. The resulting sign change corresponds to a threshold at which the excited coherent spin waves begin to decay via four-magnon scattering. At this threshold, the magnon amplitude is given by
@@
where $u_{\bm{k}}$ and $v_{\bm{k}}$ are dimensionless parameters of the Bogoliubov transformation [41] (see Supplemental Materials for details [30]), and $M_{\mathrm{s}}$ is the saturation magnetization of the magnetic film.
We calculate the spin-wave amplitude threshold corresponding to the experimental results using this expression. Table 1 summarizes the relevant experimental parameters. For YIG-A, we use the parameters listed in the left column of Table 1. The saturation magnetization is taken as $M_{\mathrm{s}}=141\,\mathrm{mT}$, as measured by SQUID (superconducting quantum interference device) magnetometry at 300 K. The anisotropy field $H_{\mathrm{ani}}$ is estimated from the difference between the effective magnetization $M_{\mathrm{eff}}=169.6\,\mathrm{mT}$ and the saturation magnetization as $H_{\mathrm{ani}}=M_{s}-M_{\mathrm{eff}}$. The decay rate $\Gamma_{\bm{k}}$ is obtained from the spin-wave lifetime, which is estimated by dividing the decay length $l_{d}=54.21\,\mathrm{\mu m}$ (determined from fitting the data at $P_{\mathrm{mw}}=11\,\mathrm{dBm}$) by the group velocity of the spin waves. The four-magnon scattering coefficient $\lambda_{\bm{k}}$ and the Bogoliubov transformation parameters $u_{\bm{k}}$ and $v_{\bm{k}}$ are calculated from analytical expressions using the experimental parameters (see Supplemental Materials for the explicit expressions [30]). Based on these parameters, the spin-wave amplitude threshold for the instability due to four-magnon scattering is calculated to be $\mu_{0}m^{\mathrm{th}}_{x,\bm{k}}=17.8\,\mathrm{mT}$.
It is important to examine how the above discussion applies to our experiment. The black dotted lines in Figs. 3(c) and (d) indicate the threshold amplitude obtained in this way. In particular, regarding Fig. 3(c), which displays the spin-wave amplitude at $x=4\,\mathrm{\mu m}$ near the stripline, the threshold for the emergence of nonlinearity and the experimental results agree satisfactorily. This result demonstrates that the nonlinear dynamics can be quantitatively visualized using NV centers. The ability to experimentally detect and quantitatively characterize the transition from the linear to the nonlinear regime, consistent with theoretical predictions, constitutes the central achievement of the present work.
@@
External magnetic field $B_{0}$ $5.17\,\mathrm{mT} 22$$2.45\,\mathrm{mT}$
@@
Anisotropy field $H_{\mathrm{ani}}$ $28.6\,\mathrm{mT}$ $15.6\,\mathrm{mT}$
Spin-wave frequency $f_{\mathrm{sw}}$ $166\,\mathrm{MHz} 22$$243.1\,\mathrm{MHz}$
@@
Four-magnon scattering coefficient $\lambda_{\bm{k}}$ -0.155 -0.173
Bogoliucdbov transformation parameter $u_{\bm{k}}$ 1.11 1.10
@@ III.5 Magnetization reduction in the nonlinear regime
Having examined the spin-wave amplitude, we next analyze the spatial distribution of the wavenumber. The observed increase in the wavenumber reflects a decrease in the static magnetization along the $y$-direction (i.e., the direction of the saturation magnetization) $M_{y,s}$ and the corresponding effective magnetization $M_{\mathrm{eff}}$ due to an increase in the spin-wave amplitude [Fig. 4(a)] [13, 14]. To calculate the effective magnetization, we utilize the dispersion relation of surface spin waves [Eq. (1)]. At each position, the effective magnetization $M_{\mathrm{eff}}(x)$ corresponding to the spin-wave frequency $f_{\mathrm{sw}}=2165.8\,\mathrm{MHz}$ is calculated from the wavenumber $k_{x}(x)$. As the dispersion relation holds for a small spin-wave amplitude, there may be deviations from Eq. (1) for large amplitudes, particularly in nonlinear regions. However, the coherent spin-wave amplitude with the excitation microwave frequency decays rapidly near the stripline in the nonlinear region, and the amplitude is small compared to the saturation magnetization across most of the field of view. Therefore, we utilize the dispersion relation of the surface spin waves in Eq. (1) to derive the static magnetization distribution approximately.
Figure 4(c) plots the resulting spatial distribution of the static magnetization $M_{y,\mathrm{s}}(x)$. We convert the effective magnetization distribution $M_{\mathrm{eff}}(x)$ into the static magnetization distribution $M_{\mathrm{y,s}}(x)$ by removing the contribution from the surface anisotropy $\quantity(M_{\mathrm{y,s}}(x)=M_{\mathrm{eff}}(x)+H_{\mathrm{ani}})$. Figure 4(c) shows that, in regions with small input microwave power, though there is a uniform reduction as the power increases, the static magnetization remains almost constant within the field of view. However, in addition to the uniform reduction, significant spatial variation within the field of view appears above $P_{\mathrm{mw}}=19\,\mathrm{dBm}$, and at the maximum input power ($P_{\mathrm{mw}}=25\,\mathrm{dBm}$), a variation of about $6\,\mathrm{mT}$ within the field of view is observed.
@@
It is instructive to examine how the surface spin waves contribute to the reduction in static magnetization [Fig. 4(b)]. When considering only the contribution from the coherent surface spin waves with the excitation frequency, the magnetization in the $y$-direction at a specific position and time is expressed as
[…]
@@
$M^{{}^{\prime}}_{y,\mathrm{s}}=\frac{1}{2\pi}\int_{0}^{2\pi}\sqrt{M^{2}_{s}-\quantity(m_{x}\cos\theta)^{2}-\quantity(\eta_{k}m_{x}\sin\theta)^{2}}\,d\theta.$ (17)
Figure 4(d) shows the effective static magnetization distribution $M^{{}^{\prime}}_{y,\mathrm{s}}(x)$, which is calculated from the spin-wave amplitude distribution in Fig. 3(a) using Eq. (17). The left edge of the plot is defined at $x=0.91\,\mathrm{\mu m}$. At small input microwave powers in the linear regime, $M^{{}^{\prime}}_{y,\mathrm{s}}$ exhibits a slight decrease only near the left edge of the field of view ($x\lesssim 10$ $\mathrm{\mu m}$), which is in clear contrast to the observation that the static magnetization derived from the wavenumber $M_{\mathrm{y,s}}(x)$ shown in Fig. 4(c)] shows a prominent decrease across the entire field of view. These results imply that the decay of the surface spin waves is not solely due to relaxation into the lattice but also involves scattering by other magnons, such as thermal magnons. In other words, the excited surface spin waves are transformed into other modes, leading to a persistent reduction in the static magnetization across the field of view [41].
Around maximum input microwave powers in the nonlinear regime, the reduction in static magnetization attributable to surface spin waves becomes almost negligible at a distance of about $5\,\mathrm{\mu m}$ from the stripline. In contrast, the static magnetization estimated from the wavenumber [Fig. 4(c)] continues to change gradually across the field of view. This observation suggests that most of the modulation in magnetization is caused by additional magnons generated through four-magnon scattering from the excited surface spin waves.
As a note, we investigate the possible role of Joule heating resulting from the application of a large-amplitude microwave, which might cause the sample to heat up. Such heating could potentially create a temperature gradient along the $x$-direction, thereby causing a spatial variation in the saturation magnetization and wavenumber. However, we can safely exclude this possibility. We perform similar measurements at a different magnetic field, where the spin-wave excitation is more inefficient and the spin waves are less likely to reach the nonlinear regime, and confirm that the wavenumber along the $x$-direction remains almost constant even at high microwave power (see Supplemental Materials for details [30]).